# Irreversibility and symmetry breaking in the creation and annihilation of defects in active living matter

### v1

@@ II Results
The polar grain boundary connected to the defect core is topologically constrained and persists throughout the defect’s lifetime. As a result, the same broken symmetry influences the defect annihilation process, as illustrated in Fig. 5b. When a $^{1}/_{2}$ defect approaches annihilation, two key forces dictate its trajectory. The first is the elastic attraction to the $^{1}/_{2}$ defect with which it is annihilating, and the second is the active self-propulsion force. The chiral torque acting on the defect causes it to rotate, so it does not self-propel directly toward the $^{1}/_{2}$ defect; however, the elastic forces prevent complete escape. Consequently, the resulting trajectory forms an inward spiral toward annihilation, as depicted in Figs. 3e-f. In contrast to the creation process, annihilation is independent of $\varphi^{+}$ (assuming $\delta=0$). Thus, defects can annihilate even when $\varphi^{+}\neq 0$, as shown in Figs. 2c-d. Additionally, the reverse of this process is equally likely, resulting in the observed “up” and “down” configurations during annihilation.
@@
Obtaining accurate statistics of the joint distribution of cellular or bacterial trajectories is challenging because of the high dimensionality of such a dataset as estimating KLD suffers from severe undersampling of the phase space Ro et al. (2022). Here, defects provide physically motivated coarse-graining of the micro-scale motion into a mesoscopically observable feature. As a consequence, all the estimates reported here are lower bounds to the actual EP.
@@
Next, we turn to analyze creation and annihilation events. Figures 2c-d show that the distribution of $^{1}/_{2}$ is different close to creation compared to annihilation. Similar broken symmetries can also be observed in the average trajectories (Figs. 3e-f) and the flow fields (Extended Data Figs. 1 and 2). To quantify this, Table 2 details the results for EP estimates. The dimension $D$ indicated the length of the trajectories used (starting at the event). Indeed, our results show positive EP both for bacteria and cells.
@@
*Note that the KLD is a non-negative quantity. However, the estimator are random variable with an error that is normally distributed with the noted standard deviation. Hence, zero EP my result in a negative estimate that is several times the standard error. In statistical terms, the hypothesis that $EP=0$ cannot be negated.
@@
Event creation annihilation cross creation annihilation cross
$EP(D=2)$ $075\pm 0.008 0.0$$4\pm 0.008 0.0$$01\pm 0.005^{*}$ $0.63\pm 0.03$ $0.55\pm 0.03$ $0.4\pm 0.3$
$EP(D=5)$ $0.11\pm 0.01$ $14\pm 0.02 0.0$$06\pm 0.02$ $0.46\pm 0.04$ $47\pm 0.05 0.0$$04\pm 0.03^{*}$
Focusing on time-reversal, running the experimental movies backward, creation will appear as annihilation and vice versa. To this end, Table 2 shows the cross entropy between forward-creation/backward-annihilation and forward-annihilation/backward-creation. The dimension $D$ indicated the length of the trajectories used (starting at the event). The positive cross-entropy indicated that creation and annihilation are not time-symmetric, both for bacteria and cells.
Comparing the total EP rates, we need to consider the rate of creation and annihilation events. For bacteria, the time between events is about 1 sec (50 frames), hence the total EP due to movement is about 50 times larger than due creation-annihilation. For cells, the time between creation and annihilation is about 2 hrs (24 frames). Hence, EP due to movement is comparable to EP due to creation-annihilation. These differences are consistent with the observation that in bacteria, $^{1}/_{2}$ defects appear self-propelled in the sense that their trajectories are ballistic over short time periods Yashunsky et al. (2024), implying a relatively large EP. With cells, both $\pm^{1}/_{2}$ defects are diffusive Blanch-Mercader et al. (2018), their movement in the bulk is close to reversible, and EP is most apparent in creation and annihilation.
@@ Conclusions
Living systems constantly produce and consume energy, keeping them far from thermodynamic equilibrium. This continuous energy input drives self-organization, often breaking spatial symmetries in ways that shape the biological function Bowick et al. (2022); Pisegna et al. (2024); Ardaševa et al. (2025). A compelling hypothesis suggests that living systems leverage topological mechanisms to regulate their organization, providing a unifying framework for understanding emergent behaviors across biological systems Shankar et al. (2022); Andersen et al. (2025).
Here, we explore this idea in two seemingly different systems: Bacillus subtilis bacterial swarms and human bronchial epithelial cells. Despite stark differences in biological function and physical properties, both systems exhibit active nematic organization, in which rod-shaped bacteria and elongated polygonal cells align and form dynamic topological defects. While bacteria propel in a thin fluid layer at low Reynolds numbers ($\sim 10^{-4}$) Be’er and Ariel (2019); Aranson (2022), HBECs move via substrate adhesion and traction forces. In both cases, motion is highly dissipative and sustained only through continuous energy consumption at the cellular level Vaidžiulytė et al. (2022). Both systems reveal nematic order, where $\pm^{1}/_{2}$ defects undergo continuing creation and annihilation events between defects pairs with opposite charge.
@@
The non-equilibrium and irreversible nature of active living systems is unequivocal. Because bacteria and cells are self propelled, the dynamics is out of equilibrium even at the level of the individual. Therefore, the theoretical prediction is that irreversibility should be observed at the single-cell level. However, the difficulty in capturing irreversibility from individual positions and orientations is reasonable. For any observed single-cell trajectory, one can imagine a cell following the reversed path. Thus, EP from single cells is only expected to be captured by adding intra-cellular degrees of freedom corresponding to chemical and biological function, or possibly also the underlying self-propulsion mechanism.
@@
We can think of defects as physically intuitive, virtual meso-scopic degrees of freedom. Virtual in the sense that they are not the particles of which the system is made of. Meso-scopic since each one is defined by a small local collective on several close cells. Our findings show that irreversibility is statistically evident locally at the level of the collective. Moreover, we find that defect creation and annihilation are not reciprocal processes that general entropy. This is also physically intuitive as continuous creation of defects in nematic systems is a hallmark of activity. In both experimental systems, entropy production during defect creation and annihilation accounts for a disproportionately large fraction of total entropy production. In cells, it is an order of magnitude higher, suggesting that these events are key drivers of non-equilibrium dynamics rather than incidental fluctuations. Still, one should remember that our estimates are merely lower bounds for the actual EP, which means that quantitative comparison is problematic.
We find that mirror symmetry, predicted by linear nematodynamic theory, spontaneously break in the process of creation and annihilation. Specifically, when close to each other, $^{1}/_{2}$ spontaneously align perpendicularly (“up” or “down”) to the pair axis (Fig. 2 e-g). Each of these configurations follows spiral-like trajectories rotating either in the clockwise (“up”) or counterclockwise (“down”) direction (Fig. 3 c-f), further indicating spontaneous breaking of the mirror symmetry.
We propose that directional self-propulsion disrupts the nematic symmetry that otherwise captures the coarse-grained behavior of bacterial swarms and epithelial cell monolayers.
This reinforces the role of topological defects in understanding principles of self-organization in living systems.
@@
Our results challenge the standard active nematic description often applied to active living matter. By uncovering these universal physical principles, we take a step toward a deeper understanding of biological self-organization, revealing the fundamental role of symmetry in shaping life at multiple scales.
@@ III Methods
Image analysis was conducted using custom Python code. To obtain the velocity field from microscopic images, we employed the Gunnar Farneback method available from OpenCV library producing displacement vectors for every pixel rather than a sparse set of points. The director field was computed using the second-moment matrix, implemented with the structure tensor function from the skimage Python library. Both the flow and director fields were computed with an averaging window size of $2.5~\mu m\times 2.5~\mu m$ for bacteria and $11.1~\mu m\times 11.1~\mu m$ for cells.
@@
Nematic defects were identified using the winding number method, which quantifies the degree of rotation in the nematic field around any closed loop of nearest neighbors. When the loop encircles a defect, the winding number reflects the charge of that defect. The orientation of the defects was calculated following the method outlined by Vromans & Giomi (2016) Vromans and Giomi (2016). Angles are discrete, at integer multiples of $\pi/50$. Defect trajectories were automatically tracked using the TrackMate plugin in ImageJ software, which was combined with defect orientation data. This generated a list of defects along with their charges, angles, and locations.
Overall, the data on a single $^{1}/_{2}$ defect trajectory is a sequence of vectors $\textbf{p}_{0},...,\textbf{p}_{K}$, where each vector pi is in $R^{3}$, i.e., each frame consists of 3 scalar values, corresponding to the xyx-y coordinates and an angle, corresponding to the orientation of the $^{1}/_{2}$ defect. The data on a single $^{1}/_{2}$ defect trajectory is a sequence of vectors $\textbf{n}_{0},...,\textbf{n}_{K}$, where each vector $\textbf{n}_{i}$ is in $R^{5}$, i.e., each frame consists of 5 scalar values, corresponding to the x-y coordinates and the three angles, corresponding to the three directions of the $^{1}/_{2}$ defect. They may not be exactly $2\pi/3$ apart. Velocities were taken as the displacements between consecutive frames with no further smoothing.
@@
Here, we provide details for the statistical estimates presented in table 3.
@@
For each $^{1}/_{2}$ defect, the detection algorithm described above provides a sequences of positions in 2D and three orientation angles, corresponding to the directions of the three legs $\theta_{1},\theta_{2},\theta_{3}\in[0,2\pi)$. We define the average defect direction (taking into account the rank three tensor symmetry) as $p=\sum_{i}(cos\theta_{i},\sin\theta_{i})$. The angle corresponding to the defect orientation is then $\varphi=\tan^{-1}(p_{y}/p_{x})/3$. Note that a function such as atan2 should be used to obtain an angle in $0,2\pi)$. Finally, the projection of defect orientation in the velocity direction (angle $\psi$) is given by $\cos 3(\varphi-\psi)$. The table shows the mean over all defects $\pm$ the standard error.
@@
First, the data for each defect trajectory or defect-pair was reduced to include only a few relative angles, as described in the results section. Looking only at the first (last) $D$ samples right after (before) a creation (annihilation) event, each event corresponds to an effective vector $\textbf{f}_{i}$. For example, looking only at the relative angle between the $^{1}/_{2}$ and $^{1}/_{2}$ defects, each creation event is reduced to a $D$-dimensional vector $\textbf{f}_{i}=(\Theta_{1}\dots\Theta_{D})$, where $\Theta_{1}$ is the angle closest to creation. The time-reversed trajectory is simply $\textbf{b}_{i}=(\Theta_{D}\dots\Theta_{1})$.
The informatic entropy production is the Kullback-Leibler divergence between the distribution of forward observations $\textbf{f}_{i}$ and backward ones $\textbf{b}_{i}$. Hence, our experimental data provides us with samples from both spaces. In order to evaluate EP, we look at the sequence of projections and find the KLD between the density of forward and backward trajectories. In backward trajectories, the sequence of positions is inverted, and the velocity changes sign. Defect angles do not change; hence the projection also changes sign. Cutting trajectories into 2-frame snippets, we obtain a coarse-grained picture, which is a lower bound of the actual EP.
Finally, cross entropy between the annihilation and creation is defined as the sum of KLD(forward creation $||$ backward annihilation) + KLD(forward annihilation $||$ backward creation ). See below for the estimation method.

### v2

@@ II Results
The polar grain boundary connected to the defect core is topologically constrained and persists throughout the defect’s lifetime. As a result, the same broken symmetry influences the defect annihilation process, as illustrated in Fig. 5b. When a $^{1}/_{2}$ defect approaches annihilation, two key forces dictate its trajectory. The first is the elastic attraction to the $^{1}/_{2}$ defect with which it is annihilating, and the second is the active self-propulsion force. The chiral torque acting on the defect causes it to rotate, so it does not self-propel directly toward the $^{1}/_{2}$ defect; however, the elastic forces prevent complete escape. Consequently, the resulting trajectory forms an inward spiral toward annihilation, as depicted in Figs. 3e-f. In contrast to the creation process, annihilation is independent of $\varphi^{+}$ (assuming $\delta=0$). Thus, defects can annihilate even when $\varphi^{+}\neq 0$, as shown in Figs. 2c-d. Additionally, the reverse of this process is equally likely, resulting in the observed “up’ and “down’ configurations during annihilation.
In this section we provide a possible physical explanation to the symmetry breaking observed above. Cellular systems such as those studied here are often analyzed using the framework of active nematics with good agreement. In particular, the constant turnover of half integer defects, in which $^{1}/_{2}$ defects are created through bend instability and self-propulsion, indicates that the system behaves as an extensile active nematic. However, in addition to nematic order, the bacteria and cells studied here also exhibit polar symmetry through their directed self-propulsion mechanism Zhou et al. (2014); Be’er and Ariel (2019); Vaidžiulytė et al. (2022); Lo Vecchio et al. (2024); Lacroix et al. (2024). Combining nematic interactions with active stresses of various symmetries can significantly alter active nematic behavior Patelli et al. (2019); Ardaševa et al. (2025); Großmann et al. (2020); de Graaf Sousa et al. (2025). In particular, breaking the polar symmetry of the nematic field at the microscopic level can lead to a break in chiral symmetry at the macroscopic level.
To describe the orientation of the cells we utilize a nematic tensor field, $\underline{\underline{Q}}$, which governs the steric interactions between the cells. In addition, we describe the head-tail symmetry of the cells by a polar vector field, $\underline{p}$, which governs the self-propulsion of the cells against the substrate. This can be visualized as two superimposed parallel fields: one with lines describing the long axes of the cells, and another with arrows for the head direction (see Fig. 5a). Both $\underline{p}$ and $\underline{p}$ are compatible with the same nematic field.
We model the system using an active nemato-polar framework in which we simultaneously simulate $\underline{\underline{Q}}$ and $\underline{p}$. The dynamics of $\underline{p}$ are constrained to follow the dynamics of $\underline{\underline{Q}}$ keeping the two fields parallel. To generate the motion of the cells we include two active terms: a standard extensile active stress, $\sigma^{a}=\zeta_{Q}\underline{\underline{Q}}$, and a body force describing cellular self-propulsion against the substrate, $\underline{f}=\zeta_{p}\underline{p}$. See SI for details.
In a typical extensile active nematic, defects are created when a bend distortion in the director field becomes unstable. The bend distortions are characterized by a non-zero curl, which is unsigned for the nematic field. However, in the polar field the bend distortion is signed and therefore breaks mirror symmetry. The resulting half integer defects are incompatible with the polar field. This leads to a grain boundary that connects the two defects, across which the polar field changes sign, see blue line on Fig. 5a. Notably, the mirror image of this process produces defects with opposite chirality and is equally likely, indicating that defect chirality selection is driven by spontaneous symmetry breaking during creation. In this context, distinct “up” and “down” defect configurations can emerge during creation.
We simulate the defect creation process by initializing the fields with a single prominent bend distortion at the center of the simulation domain. Upon simulation, the bend distortion nucleates a pair of defects that separate and follow curved trajectories, see Fig. 5a (inset), recreating those observed in experiments, see Fig. 3e-f. We simulate the annihilation process by initializing the simulation with a pair of defects such that $\varphi^{+}=\pm\pi$, (see Fig. 5b), as expected in active nematic. Simulated annihilations also show curved trajectories (Fig. 5b inset), resembling the experimental observations (Fig. 3e-f). We quantify the co-rotation of the defect pair by plotting the total change in $\gamma$ throughout the creation or annihilation process, see Fig. 5c. We see that the degree of co-rotation increases with the strength of the polar forcing, $\zeta_{p}$, and is larger during annihilation. We also observe rotation of the $^{1}/_{2}$ defect, quantified by $\Delta\varphi^{+}$ at the end of the simulation, shown in Fig. 5d. The rotation of the defect increases with the magnitude of the polar force. Consistent with our experimental observations (Fig. 3e–f), the rotation is more pronounced during annihilation than during creation. Importantly, the chirality of the simulations can be reversed either by mirroring the fields or, equivalently, by changing $\zeta_{p}\rightarrow-\zeta_{p}$. This procedure allows to simulate the system in either the “up” or “down” configuration, shown as green or yellow in Figs. 5e-f. If we remove the active body force, $\zeta_{p}=0$, we recover the straight trajectories expected for typical active nematics.
In our experiments, the flow profile around a $^{1}/_{2}$ defect is offset relative to the symmetry axis of the defect, see Fig. 4. We can analytically calculate how the introduction of the polar body force influences the flow around a $^{1}/_{2}$ defect and observe a similar offset, see Fig. 5e and SI for details. The polar body force breaks the mirror symmetry of the defect, resulting in a central flux at an angle $\phi_{V}$ relative to the tail of the $^{1}/_{2}$ defect. The magnitude of this angular offset depends on the angular position of the grain boundary ($\phi_{g}$) and the scale of the polar active force, see Fig. 5f. The distribution of $\phi_{V}$ measured from our experiments has a peak at $\approx\pm\pi/4$ for bacteria and $\approx\pm\pi/3$ for cells, see Fig. 4 and Extended Fig. 6a-b. Such an offset would indicate that the polar forcing accounts for at least $25\%$ of the flow around the defect. The polar body force also introduces a net vorticity around the core of the defect, see Fig. 5g. While the net flux at the core of the defect causes it to self-propel, the net vorticity induces a torque that causes defects to rotate. Similar to the velocity offset, the magnitude of the rotation is controlled by the magnitude of the polar active force, but the direction of rotation depends only on the sign of $\zeta_{p}$, hence the chirality of the defect.
Taken together, these observations explain the spontaneous symmetry breaking into “up” and “down” configurations observed in experiments. Defects are generated through a bend instability due to the extensile active stress, hence with $\varphi^{+}\approx 0$. This process breaks the mirror symmetry of the polar field and introduces a grain boundary across which $\underline{p}$ changes sign (blue dashed line in Fig. 5a). The grain boundary connects the defects and results in a net curl around the $^{1}/_{2}$ defect which translates to a net torque (green arrow in Fig. 5a). This torque causes the $^{1}/_{2}$ defect to rotate as it self-propels, thus $\dot{\varphi}^{+}\neq 0$. Combined with the defect self-propulsion, this rotation results in the defect taking a curved trajectory, spiraling away from its $^{1}/_{2}$ equivalent (Fig. 5a). The sign of the torque depends on the sign of the curl of the initial bend distortion. Thus, it arises from a spontaneous break in symmetry. A counter clockwise net torque on the defect leads to $\dot{\varphi}^{+}>0$ and we observe the “up” configuration, as shown in Fig. 5a.
The polar grain boundary connected to the defect core is topologically constrained and persists throughout the defect’s lifetime. Thus, a $^{1}/_{2}$ defect has an offset flow profile throughout with a chirality that is determined at creation (Fig. 5e). The latter drives the bimodal distribution of $^{1}/_{2}$ defect velocities as “up” and “down” defects self-propel with an angular offset (Figs. 45f, and Extended Fig. 6. This indicates that the chiral symmetry of a $^{1}/_{2}$ is broken at creation and persists until annihilation, where it manifests again in that process.
During annihilation, the $^{1}/_{2}$ defect tail initially points away from the $^{1}/_{2}$ defect, hence $\varphi^{+}=\pm\pi$. In this case a clockwise net torque around the $^{1}/_{2}$ defect results in $\dot{\varphi}^{+}<0$ which causes the defect to rotate into the “up” annihilation configuration $\varphi^{+}>0$ (Fig. 5b). The latter is the chirality associated with defects that originated in the “down” configuration at creation. Hence, we expect the chirality of $^{1}/_{2}$ defects created in the “up” configuration to be identical to those that annihilate in the “down” configuration, as indeed observed in experiments (Fig. 2). As in creation, this rotation, combines with defect self-propulsion, results in a curved trajectory (Fig. 5d). In contrast to the creation process, annihilation is independent of $\varphi^{+}$ (assuming $\delta=0$). Thus, defects can annihilate even when $\varphi^{+}\neq 0$, as shown in Figs. 2c-d.
@@
Obtaining accurate statistics of the joint distribution of cellular or bacterial trajectories is challenging because of the high dimensionality of such a dataset as estimating KLD suffers from severe undersampling of the phase space Ro et al. (2022). Here, defects provide physically motivated coarse-graining of the micro-scale motion into a meso-scopically observable feature. As a consequence, all the estimates reported here are lower bounds to the actual EP.
@@
Next, we turn to analyze creation and annihilation events. Figures 2c-d show that the distribution of $\varphi^{+}$ is different close to creation compared to annihilation. Similar broken symmetries can also be observed in the average trajectories (Figs. 3e-f) and the flow fields (Extended Data Figs. 1 and 2).
Focusing on time-reversal, running the experimental movies backward, creation will appear as annihilation and vice versa. To this end, Table 2, column labeled cross, shows the symmetric KL divergence (Jeffreys divergence) between forward-creation/backward-annihilation and forward-annihilation/backward-creation,
${\color[rgb]{0,0,0}\frac{1}{2}D_{\rm KL}(\overrightarrow{\rm creation}||}$ ${\color[rgb]{0,0,0}\overleftarrow{\rm annihilation})+}$ (1)
${\color[rgb]{0,0,0}\frac{1}{2}D_{\rm KL}(\overrightarrow{\rm annihilation}||\overleftarrow{\rm creation})}.$
The positive symmetric divergence indicated that creation and annihilation are not time-symmetric, both for bacteria and cells.
Comparing the total EP rates, we need to consider the rate of creation and annihilation events. For bacteria, the time between events is about 1 sec (50 frames), hence the total EP due to movement is about 50 times larger than due creation-annihilation. For cells, the time between creation and annihilation is about 2 hrs (24 frames). Hence, EP due to movement is comparable to EP due to creation-annihilation. These differences are consistent with the observation that in bacteria, $^{1}/_{2}$ defects appear self-propelled in the sense that their trajectories are ballistic over short time periods Yashunsky et al. (2024), implying a relatively large EP. With cells, both $\pm^{1}/_{2}$ defects are diffusive (or the timescale of ballistic motion is relatively much shorter) Blanch-Mercader et al. (2018), their movement in the bulk is close to reversible, and EP is most apparent in creation and annihilation.
@@
$0.05\pm 0.01$ $0.5\pm 0.2$
@@ Conclusions
By studying defect dynamics in two distinct active living systems—swarming Bacillus subtilis bacteria and human bronchial epithelial cell (HBEC) monolayers—we demonstrate that topological defect creation and annihilation are fundamental, symmetry-breaking processes that drive these systems far from equilibrium. Despite orders-of-magnitude differences in timescales and propulsion mechanisms, both systems exhibit nematic half-integer defects whose pair dynamics break both mirror symmetry and time-reversal symmetry. Moreover, $^{1}/_{2}$ defects spontaneously adopt persistent “up” or “down” chiral configurations and follow spiral trajectories.
Irreversibility and Entropy Production
@@
The non-equilibrium and irreversible nature of active living systems is unequivocal. Because bacteria and cells are self-propelled, the dynamics is out of equilibrium even at the level of the individual. Therefore, the theoretical prediction is that irreversibility should be observed at the single-cell level. However, the difficulty in capturing irreversibility from individual trajectories is intuitively clear. For any observed single-cell trajectory, one can imagine a cell tracing the reversed path. Thus, EP from single cells is only expected to be captured by adding intra-cellular degrees of freedom corresponding to chemical and biological function or the underlying self-propulsion mechanism. From this respect, defects are physically intuitive, virtual meso-scopic degrees of freedom. Virtual in the sense that they are not the particles of which the system is made of. Meso-scopic since each one is defined by a small local collective of several close cells.
@@
Our findings thus show that irreversibility is statistically evident locally at the level of the collective; see also Tan et al. (2019) for measurements of topological entropy and Radhakrishnan et al. (2025) for a theoretical study. Moreover, we find that defect creation and annihilation are not reciprocal processes and that they generate entropy. This is also physically intuitive as the continuous creation of defects in nematic systems is a hallmark of activity. In both experimental systems, entropy production during defect creation and annihilation accounts for a disproportionately large fraction of the total entropy production. In cells, it is an order of magnitude higher, suggesting that these events are key drivers of non-equilibrium dynamics rather than incidental fluctuations. Still, one should remember that our estimates are merely lower bounds for the actual EP, which means that quantitative comparison is still lacking.
Overall, entropy production estimates bring forth some fundamental differences between swarming bacteria and crawling cells regarding the main sources of entropy production (defect motion in bacteria and creation-annihilation in cells). This points to the physical mechanisms underlying departure from equilibrium.
Spontaneous Symmetry Breaking and the Polar-Nematic Model
Our experiments reveal that topological defect creation and annihilation in living active nematics spontaneously break mirror symmetry—a phenomenon absent in standard active nematic theory. Specifically, $^{1}/_{2}$ defects near a $^{1}/_{2}$ defect adopt one of two distinct chiral configurations, exhibiting bimodal orientation distributions and persistent spiral trajectories. This chiral symmetry breaking is established during defect creation and persists until annihilation.
To explain these observations, we introduce an active nemato-polar model that couples a polar vector field $\bm{p}$—representing head–tail asymmetric self-propulsion—to the nematic tensor order parameter $\bm{Q}$. A half-integer topological defect in the nematic field imposes a $\pi$-discontinuity (grain boundary) in the polar field, breaking the mirror symmetry of the $^{1}/_{2}$ defect core. This generates a net active torque that rotates the defect as it self-propels. The torque sign, selected spontaneously during nucleation, determines the chirality and drives the spiral motion of defect pairs.
The model reproduces key experimental signatures: the bimodal distribution of orientation angle $\varphi^{+}$, the chiral rotation quantified by $\Delta\gamma$, and the flow field offset around $^{1}/_{2}$ defects. It also explains the stark contrast with achiral active nematics (e.g., microtubule–kinesin suspensions), where the absence of a polar field yields dynamics consistent with standard nematic theory.
We suggest that spontaneous symmetry breaking in living active matter stems directly from the intrinsic polar activity of its constituents. A complete description therefore requires accounting for the interplay between nematic alignment and polar self-propulsion, revealing how microscopic asymmetry is imprinted onto meso-scopic topological degrees of freedom.
@@
Overall, our results challenge the standard active nematic description often applied to living active matter. By uncovering these universal physical principles, we advance understanding of biological self-organization, revealing the fundamental role of symmetry in shaping life at multiple scales.
@@ III Methods
Image analysis was conducted using custom Python code. To obtain the velocity field from microscopic images, we employed the Gunnar Farneback method available from OpenCV library producing displacement vectors for every pixel rather than a sparse set of points Farnebäck (2003). The director field was computed using the second-moment matrix, implemented with the structure tensor function from the skimage Python library. Both the flow and director fields were computed with an averaging window size of $2.5~\mu m\times 2.5~\mu m$ for bacteria and $11.1~\mu m\times 11.1~\mu m$ for cells.
@@
Topological defects were identified by evaluating the winding of the director field. We computed the topological charge $k$ at each grid point by summing the angular differences of the director along a closed loop around a $2\times 2$ unit cell. Because the nematic phase is apolar, all angular differences were wrapped to the interval $(-\pi/2,\pi/2]$. Grid points with a net winding of $\pm\pi$ were labeled as defects with strength $k=\pm^{1}/_{2}$.
Defect orientations were determined using a direct interpolation method. The local director field was interpolated onto a circular path centered on each defect. We then extracted the orientation angle by maximizing the agreement between this interpolated profile and the analytical director field for a defect of strength $k$. For $^{1}/_{2}$ defects, this yielded a single orientation axis that points toward the comet head. For $^{1}/_{2}$ defects, the method resolved the three-fold symmetry of the core and returned the three equivalent orientation angles separated by $120^{\circ}$.
Defect trajectories were tracked using the trackpy Python library Crocker and Grier (1996). The tracking results were merged with the orientation data to produce a list of defects with their charges, orientations, and positions. Overall, the data on a single $^{1}/_{2}$ defect trajectory is a sequence of vectors $\textbf{p}_{0},...,\textbf{p}_{K}$, where each vector pi is in $R^{3}$, i.e., each frame consists of 3 scalar values, corresponding to the xyx-y coordinates and an angle, corresponding to the orientation of the $^{1}/_{2}$ defect. The data on a single $^{1}/_{2}$ defect trajectory is a sequence of vectors $\textbf{n}_{0},...,\textbf{n}_{K}$, where each vector $\textbf{n}_{i}$ is in $R^{5}$, i.e., each frame consists of 5 scalar values, corresponding to the x-y coordinates and the three angles, corresponding to the three directions of the $^{1}/_{2}$ defect. They may not be exactly $2\pi/3$ apart. Velocities were taken as the displacements between consecutive frames with no further smoothing.
@@
Here, we provide details for the statistical estimates presented in table  3.
@@
For each $^{1}/_{2}$ defect, the detection algorithm described above provides a sequences of positions in 2D and three orientation angles, corresponding to the directions of the three legs $\theta_{1},\theta_{2},\theta_{3}\in[0,2\pi)$. We define the average defect direction (taking into account the rank three tensor symmetry) as $p=\sum_{i}(cos3\theta_{i},\sin 3\theta_{i})$. The angle corresponding to the defect orientation is then $\varphi=\tan^{-1}(p_{y}/p_{x})/3$. Note that a function such as atan2 should be used to obtain an angle in $0,2\pi)$. Finally, the projection of defect orientation in the velocity direction (angle $\psi$) is given by $\cos 3(\varphi-\psi)$. The table shows the mean over all defects $\pm$ the standard error.
The informatic entropy production is the Kullback-Leibler divergence between the distribution of forward observations $\textbf{f}_{i}$ and backward ones $\textbf{b}_{i}$. Hence, our experimental data provides us with samples from both spaces. In backward trajectories, the sequence of positions is inverted, and the velocity changes sign (because velocities are a forward difference rule). However, defect angles do not change; hence the projection also changes sign, hence the distribution of forward projections and backward ones differ. Cutting trajectories into 2-frame snippets, we obtain a coarse-grained picture, which is a lower bound of the actual EP.
@@
There are 4 angles associated with the creation/annihilation processes: there related to the defect orientation - $\varphi^{+}$, $\varphi^{-}$ and $\delta$ - and one related to the line connecting the defect centers, $\gamma$. We have defined $\gamma$ relative to the angle between the $\pm^{1}/_{2}$ defects at the time of the event. Hence, it is not a state function (it cannot be obtained from a single frame). Hence, using it may introduce irreversibility, because its measurement is with respect to a fixed event which is either set in the past or future. Similarly, $\varphi^{+}$ and $\varphi^{-}$ are relative to $\gamma$. However, both $\delta$ and $\Delta\varphi=\varphi^{+}-\varphi^{-}$ are not (here, $\varphi^{-}$ is the angle of the leg which is closes to the comet-tail of the +). Therefore, one should use these angles when looking for irreversibility.
At each event, creation or annihilation, $\delta$ approaches zero. Hence, this sense, creation and annihilation are time-symmetric. As we demonstrate below, the statistics of how $\delta\to 0$ may be more subtle. However, substituting $\delta=0$, $\Delta\varphi=2\varphi^{+}/3+\pi/3$, which is not expected to be reversible.
More specifically, for each event we look at a short sequence of $D$ snapshots (before annihilation or after creation). Looking at $\delta$ and $\Delta\varphi$, a forward trajectory yield a $2D$ dimensional vector $\textbf{f}_{i}=(\delta_{1},\Delta\varphi_{1},\dots,\delta_{D},\Delta\varphi_{D})$. The time-reversed trajectory is then $\textbf{b}_{i}=(\delta_{D},\Delta\varphi_{D},\dots,\delta_{1},\Delta\varphi_{1})$. Note that it is imperative that the order of $\delta$ and $\Delta\varphi$ does not change.
Finally, the symmetric KL divergence (Jeffreys divergence) between the annihilation and creation is defined as half the sum of KLD(forward creation $||$ backward annihilation) + KLD(forward annihilation $||$ backward creation ). See below for the estimation method.
The results may depend on the choice of $D$. On the one hand, smaller $D$ implies more coarse graining, resulting in a looser lower bound. On the other hand, higher $D$ implies a higher dimension for the estimate of the divergence, usually leading to larger statistical errors. Since, for any $D$, the estimate provides a lower bound for the actual EP, We may optimize its value. We use $D=4$ for bacteria and $D=11$ for cells.